40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 2 00:54:08 2017, Koji, Update, ASS, Xarm ASS restoration work 7x
    Reply  Fri Jun 2 01:22:50 2017, gautam, Update, ASS, ASS restoration work ASS_Y_recovery.pngASS_Repairs.zipOLs.png
       Reply  Wed Jun 7 14:11:49 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation CoilTFs.pdfY_OL_calib_check.png
          Reply  Wed Jun 7 17:45:11 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation CoilTFs.pdf
             Reply  Thu Jun 8 02:11:28 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation AnalogCheck.pdfAnalogCheck_zoom.pdfArmCtrl.pdf
                Reply  Mon Jun 12 17:42:39 2017, gautam, Update, ASS, ETMY Oplev Pentek board pulled out 
                   Reply  Wed Jun 14 18:15:06 2017, gautam, Update, ASS, ETMY Oplev restored 
                      Reply  Thu Jun 15 01:56:50 2017, gautam, Update, ASS, ETMY Oplev restored ITMY_OLspec.pdf
                         Reply  Tue Jun 27 16:11:28 2017, gautam, Update, Electronics, Coil whitening ETMYT_1182669422.mp4ITMY_whitening_studies.pdf
Message ID: 13048     Entry time: Wed Jun 7 14:11:49 2017     In reply to: 13033     Reply to this: 13051
Author: gautam 
Type: Update 
Category: ASS 
Subject: Y-arm coil driver electronics investigation 

Rana suggested taking a look at the Y-arm test mass actuator TFs (measured by driving the coils one at a time, with only local damping loops on, using the Oplev to measure the response to a given drive). Attached are the results from this measurement (I used the Oplev pitch error signal for all 8 measurements). Although the magnitude response for all coils have the expected 1/f^2 shape, there seems to be some significant (~10dB) asymmetry in both the ETM and ITM coils. The phase-response is also not well understood. If we are just measuring the TF of a pendulum with 1 Hz resonant frequency, then at and above 10Hz, I would expect the phase to be either 0 or 180 deg. Looks like there is a notch at 60 Hz somewhere, but it is unclear to me where the ~90 degree phase at ~100Hz is coming from.

For the ITM, the UL OSEM was replaced during the 2016 summer vent - the coil that is in there is now of the short OSEM variety, perhaps it has a different number of turns or something. I don't recall any coil balancing being done after this OSEM swap. For the ETM, it is unclear to me how long this situation has been like this.

Yesterday night, I tried to measure the ASS output matrix by stepping the ITM, ETM and TTs in PIT and YAW, and looking at the response in the various ASS error signals. During this test, I found the ETM and ITM pitch and yaw error signals to be highly coupled (the input matrix was diagonal). As Rana suggested, I think the whole coil driver signal chain from DAC output to coil driver board output has to be checked before attempting to fix ASS. Results from this investigation to follow.

Note: The OSEM calibration hasn't been done in a while (though the HeNes have been swapped out), but as Attachment #2 shows, if we believe the shadow sensor calibration, then the relative calibrations of the ITM and ETM Oplevs agree. So we can directly compare the TFs for the ITM and ETM.

 

Attachment 1: CoilTFs.pdf  24 kB  | Hide | Hide all
CoilTFs.pdf
Attachment 2: Y_OL_calib_check.png  25 kB  Uploaded Wed Jun 7 15:12:45 2017  | Hide | Hide all
Y_OL_calib_check.png
ELOG V3.1.3-